OpenStudy (anonymous):

What is y=10(square root 3)to the 14th power?

OpenStudy (aum):

\[y = (10\sqrt 3)^{14} = 10^{14} * (3^{1/2})^{14} = 10^{14} * 3^{7} = 2187 \times 10^{14} = 2.187 \times 10^{17}\]

OpenStudy (aum):

The above answer is in scientific notation.